# Localized Distributional Robustness in Submodular Multi-Task Subset Selection

### v2

@@ Abstract
In this work, we approach the problem of multi-task submodular optimization with the perspective of local distributional robustness, within the neighborhood of a reference distribution which assigns an importance score to each task. We initially propose to introduce a regularization term which makes use of the relative entropy to the standard multi-task objective. We then demonstrate through duality that this novel formulation itself is equivalent to the maximization of a monotone increasing function composed with a submodular function, which may be efficiently carried out through standard greedy selection methods. This approach bridges the existing gap in the optimization of performance-robustness trade-offs in multi-task subset selection. To numerically validate our theoretical results, we test the proposed method in two different settings, one on the selection of satellites in low Earth orbit constellations in the context of a sensor selection problem involving weak-submodular functions, and the other on an image summarization task using neural networks involving submodular functions. Our method is compared with two other algorithms focused on optimizing the performance of the worst-case task, and on directly optimizing the performance on the reference distribution itself. We conclude that our novel formulation produces a solution that is locally distributional robust, and computationally inexpensive.
Keywords: Optimization, robustness, submodularity.
@@ 1 Introduction
Submodular functions have for long been the focus of extensive studies[2, 3, 4, 5], thanks to their propensity to occur naturally in many distinct domains such as economics and algorithmic game theory [6, 7, 8] and machine learning[9, 10, 11, 12, 13, 14, 15, 16, 17, 18], within many distinct problems, ranging from sensor selection [19, 5, 20, 21] to social network modeling [22]. In simplest terms, a submodular function is a set function that possesses the diminishing marginal gains property [23]. This property, at first glance, enables one to draw an analogy between submodular functions and concave functions in the continuous domain. However, interestingly, a similar analogy may be drawn between submodular functions and convex functions in the continuous domain, as well, through the use of certain extensions of submodular functions onto the continuous domain [24, 5]. The similarity of submodular functions to both convex and concave functions, the white whales of optimization literature, has also garnered interest in the study of the optimization of submodular functions.
@@
The quintessential submodular optimization problem is that of the maximization of a submodular function under a cardinality constraint. This problem entails selecting a subset out of a ground set $N$ of elements, such that the size of the selection does not exceed a preset constraint value, and the evaluation of the selection under the submodular function is as high as possible. The evaluation $f(S)$ of a subset $S\subseteq N$ by $f$ is commonly called the score or utility of the set $S$. The following formulation, where, $f$ is a submodular function, $N$ is the ground set and $K$ is a positive integer representing the cardinality constraint, expresses this problem in mathematical terms[6]:
$\begin{split}&\max_{S\subseteq N}f(S)\\ &\text{s. t. }\lvert S\rvert\leq K.\end{split}$ (1)
@@
The notion of robustness occurs very frequently in optimization literature and is the object of much consideration. Although there are differing definitions of robustness depending on the context at hand, such as being immune to removals from the solution [25], or to slight changes in parameters or input data[26], and others[27], it always encompasses the general idea of a produced solution to a problem (the output) staying correct and valid under perturbations to the input. In the present context of submodular optimization, one conception of robustness that is commonly encountered is multi-task robustness. As the name implies, with this conception, the aim is to algorithmically produce a solution set whose utility with respect to multiple submodular functions $f^{1},\ldots,f^{n}$ is satisfactory.
@@
As it stands, the notion of being “satisfactory” is intentionally left to be vague, and the problem can be formulated in multiple ways to fit the description. A straightforward interpretation with this goal in mind might lead to what we will call the worst-case formulation [5, 28, 29, 30], which aims to focus on producing a solution that maximizes the worst-performing objective function among all. Using the shorthand notation $[n]\vcentcolon=\{1,\ldots,n\}$, this formulation is as follows:
$\begin{split}&\max_{S\subseteq N}\min_{i\in[n]}f^{i}(S)\\ &\text{s. t. }\lvert S\rvert\leq K.\end{split}$ (2)
@@
A shortcoming of this formulation is the fact that it can be viewed as being too pessimistic[31], dedicating all resources to the maximization of the worst-performing objective function, at the expense of disregarding the others. In scenarios where one objective function is a clear outlier, as in it always scores significantly lower than all others, this approach may effectively lead to getting low utility on all functions in pursuit of the hopeless aim of maximizing that outlier.
@@
Another way to formulate the multi-task robust problem could be to adopt what we will call the average-case formulation [31, 32], which aims to directly optimize the arithmetic average of all the objective functions. In this case, the formulation would be
@@
$\begin{split}&\max_{S\subseteq N}\frac{1}{n}\sum_{i=1}^{n}f^{i}(S)\\ &\text{s. t. }\lvert S\rvert\leq K.\end{split}$ (3)
@@
The apparent shortcoming of this formulation is the fact that contrary to the previous formulation, it provides no guarantees whatsoever about how low the utility with respect to any individual objective function may be. In essence, one or multiple objective functions could be scoring arbitrarily badly, as long as the other objective functions are scoring well enough to compensate for the underperformers.
@@
The previous two formulations, even with their stated shortcomings, may very well be acceptable approaches, especially in the absence of any additional information. However, we argue that if we possess additional information on the nature of the objective functions, or a system at hand, for whose various tasks we use the multiple objective functions to model, we can progress to a more meaningful formulation. Suppose, for instance, that we have access to a reference distribution, an $n-$dimensional discrete probability distribution $Q\in\Delta_{n}$. In a practical setting, the reference distribution $Q$ may be obtained by a decision-maker subjectively assigning importance scores to each objective function, or by a frequentist approach, where the value assigned to each function would be given by how frequently the task that is modeled by said function is performed by the system. It is reasonable to argue that the reference distribution can facilitate incorporating the intent and priority of the main stakeholders in the problem. The incorporation of this additional information to the previous formulation leads to a generalization of Problem (3), where the reference distribution is simply assumed to be uniform. This idea is encapsulated in the following formulation:
$\begin{split}&\max_{S\subseteq N}\sum_{i=1}^{n}Q_{i}f^{i}(S)\\ &\text{s. t. }\lvert S\rvert\leq K.\end{split}$ (4)
@@
We note that this previous formulation is a generalization of Problem (2), such that, when $Q$ is a one-hot discrete distribution indicating the worst-performing objective function, this problem becomes equivalent to Problem (2).
@@
Fig. 1: Illustration of the three discrete distributions on the $3-$dimensional simplex[33] that the three discussed approaches optimize. $P_{\text{worst}}$ corresponds to the global worst-case task scenario, assigning a weight of $1$ to the worst-case task and a weight of 00 to all the others, residing on a vertex of the simplex. $P_{\text{avg}}$ assigns uniform weight to all tasks, and lies in the center of the simplex. $Q$ is the reference distribution within a neighborhood of which we want to achieve local robustness. $P^{\ast}$ is the local worst-case distribution within that neighborhood of $Q$.
@@ 1.1 Contribution
Based on this previous formulation, we present our novel robust formulation, which not only makes use of the reference distribution to simply weigh each objective function but to act as the center of a neighborhood of robustness, within which we will aim to be robust. The initial candidate formulation for this is as follows:
$\begin{split}&\max_{S\subseteq N}\min_{P\in\Delta_{n}}\sum^{n}_{i=1}P_{i}f^{i}(S)\\ &\text{s. t. }\lvert S\rvert\leq K,\;D(P\;\|\;Q)\leq R,\end{split}$ (5)
where the constraint $D(P\;\|\;Q)\leq R$ designates the requirement that we localize ourselves to a specified neighborhood of robustness, with radius $R$, only within which we will maximize our objectives, weighed by the worst possible distribution. $D(P\;\|\;Q)$ is intentionally unspecified and could be any metric or divergence. For an investigation of different statistical distances for $D(P\;\|\;Q)$, we refer the reader to Section 3.
Before we proceed with the theoretical analysis of Problem 5, we propose one last change to the formulation to make it more tractable. We relax the formulation by introducing the constraint $D(P\;\|\;Q)\leq R$ into the objective via a penalizing regularization constant $\lambda>0$, instead of strictly enforcing it. This change reflects the standard idea of introducing Lagrange multipliers, utilized frequently in continuous optimization[34]. With this final addition, our novel formulation becomes
$\begin{split}&\max_{S\subseteq N}\min_{P\in\Delta_{n}}\sum^{n}_{i=1}P_{i}f^{i}(S)+\lambda D(P\;\|\;Q)\\ &\text{s. t. }\lvert S\rvert\leq K.\end{split}$ (6)
We propose, practically, that this final formulation will be useful in many kinds of multi-objective applications, where we aim to maintain a performance threshold in all tasks involved, especially if we are in the knowledge of a reference distribution indicating the relative importance scores of each task. In particular, we will be undertaking two such applications in Section 6. The first will involve a low-earth orbit satellite constellation performing multiple tasks involving atmospheric readings and ground coverage, where in critical missions it may be of importance to have satisfactory performance in all objectives. The second application involves a task of image summarization, where each objective will represent the cosine distance of each image in a dataset to the current selection made from the dataset, intended to be its summary. It is clear that in such a task, one would expect a certain level of performance to be met with respect to all of the images, so that the resulting summarization may be considered a good one.
@@ 1.2 Related Works and Significance
Robust submodular optimization. Reference [19] is a fundamental work in establishing the notion of robustness in submodular optimization that is adopted in the current manuscript, and proposing an algorithmic solution to the worst-case formulation of Problem (2), namely, the Submodular Saturation Algorithm (SSA). This algorithm will be discussed considerably in the following sections and will constitute one of the baselines for our work to be compared against. References [35, 30] similarly adopt the same formulation of Problem (2) in consideration of robustness. The former of these works generalizes the results of [19], which deals with the cardinality-constrained problem, to cases where knapsack or other matroid constraints are present, including the case where the constraints themselves are submodular, as well. The latter shifts the focus to achieving higher computational efficiency, aiming for a fast and practical algorithm with asymptotic approximation guarantees when $n$, the number of objective functions is increased arbitrarily. Reference [31] proposes a novel notion of robustness that yet pertains to the context of multiple objective functions. The authors firstly motivate their novel approach by drawing attention to the pessimistic nature of Problem (2). They instead propose an approach that entails the maximization of the $p$th quantile of the objective functions, where $10\leq p<$ 1. To simplify, this approach may be viewed as forgoing the consideration of the worst-performing $p$th of the objective functions in favor of maximizing the remaining. Different from all of the previously mentioned works, the present work shifts the focus to the scope of distributional robustness, focusing on localizing our goal of robustness within a subset of the $n-$dimensional simplex, in a neighborhood of a reference distribution. References [25, 36] adopt one of the aforementioned different notions of robustness within the context of submodular optimization. In these works, the notion of robustness is against the post-processing of solutions for the arbitrary or adversarial deletion of a certain number of elements, rather than robustness against multiple submodular objective functions.
@@
Distributionally-robust optimization. Our work is closely related to and motivated by the distributionally robust optimization paradigm. In a very general sense, distributionally robust optimization is a paradigm where some uncertainty about the nature of the problem is governed by a probability distribution and the goal is to leverage the information we have about this governing distribution with the goal of optimization or optimal decision-making. Reference [37] approaches a convex optimization problem in the context of distributionally robust optimization. Reference [38] broadens the scope of consideration of online optimization of nonconvex functions. In particular, the authors use relative entropy regularization, which we also employ in this work. Reference [32], much like our work, deals with distributionally-robust optimization of submodular functions. The authors here work in the presence of stochastic submodular objective functions that are drawn from a given reference distribution and aim to maximize the expected value of such functions. They propose an approach that entails a variance-regularized objective, making use of the multilinear extension of submodular functions and the Momentum Frank-Wolfe algorithm. In contrast to this line of work, we leverage distributionally robust optimization for discrete multi-task subset selection and demonstrate the possibility of formulating the distributionally robust problem as a simple submodular maximization problem, essentially allowing us to produce a distributionally robust solution at no additional cost to produce a solution to the naive formulation of Problem (4).
@@ 2 Preliminaries and Background
We begin by introducing the concept of the marginal gain in set functions because it is key in defining submodular functions.
Definition 1 (Marginal gain).
Given a set function $f\vcentcolon 2^{N}\rightarrow\mathbb{R}$ and $A,B\subseteq N$, we denote $f(B\cup A)-f(B)$, the marginal gain in $f$ due to adding $A$ to $B$, by $\Delta_{f}(A|B)$.
When the set $A$ is a singleton, i.e., $A=\{a\}$, we drop the curly brackets to adopt the short-hand notation $\Delta_{f}(a|B)$.
@@
A set function $f\vcentcolon 2^{N}\rightarrow\mathbb{R}$ is submodular if for every $A\subseteq B\subseteq N$ and $e\in N\setminus B,$ it holds that
$\Delta_{f}(e|A)\geq\Delta_{f}(e|B).$ (7)
This definition of submodularity highlights the renowned diminishing marginal gains property of submodular functions, which is commonly referred to when making the analogy to discrete concave functions. The common notion in thinking about a submodular function $f$ is that it scores each set by its utility, as such, we will commonly refer to the evaluation $f(S)$ as the utility, performance, or score of set $S$.
We further introduce two additional useful properties of set functions, that are combined with submodularity in the derivation of theoretical guarantees for approximation algorithms that optimize submodular functions.
@@
A set function $f\vcentcolon 2^{N}\rightarrow\mathbb{R}$ is normalized if $f(\emptyset)=0$.
@@
A set function $f\vcentcolon 2^{N}\rightarrow\mathbb{R}$ is monotone nondecreasing if for every $A\subseteq B\subseteq N,$ we have $f(A)\leq f(B).$
Note that a set function $f$ having both of the last two properties ensures that $f\geq 0$.
In practical applications, it is often the case that the objective functions involved are not submodular, but weak-submodular [39]. In literature, there are multiple notions of weak submodularity, but for the purposes of this paper, we will be adopting one that is well-established [40, 41, 42, 39, 43], and well extends the key concept of submodular functions, the concept of diminishing marginal returns. To facilitate the definition however, we first define the weak-submodularity constant (WSC):
@@
$w_{f}\vcentcolon=\max_{(S,T,e)\in\tilde{N}}\dfrac{\Delta_{f}(e|T)}{\Delta_{f}(e|S)},$ (8)
where $\tilde{N}=\{(S,T,e)|S\subseteq T\subset N,e\in N\setminus T\}.$
An intuitive understanding of the WSC is that it is a measure of how badly the diminishing marginal gains property required for submodularity is violated. Evidently, a monotone nondecreasing set function $f$ is submodular if and only if its WSC satisfies $w_{f}\leq 1$.
With this definition in hand, we now define weak submodularity.
@@
A set function $f\vcentcolon 2^{N}\to\mathbb{R}$ is weak-submodular if it is monotone nondecreasing and its WSC satisfies $w_{f}<\infty$.

### v3

@@ Abstract
In this work, we treat the problem of multi-task submodular optimization from the perspective of local distributional robustness within the neighborhood of a reference distribution which assigns an importance score to each task. We initially propose to introduce a relative-entropy regularization term to the standard multi-task objective. We then demonstrate through duality that this novel formulation itself is equivalent to the maximization of a monotone increasing function composed with a submodular function, which may be efficiently carried out through standard greedy selection methods. This approach bridges the existing gap in the optimization of performance-robustness trade-offs in multi-task subset selection. To numerically validate our theoretical results, we test the proposed method in two different settings, one on the selection of satellites in low Earth orbit constellations in the context of a sensor selection problem involving weak-submodular functions, and the other on an image summarization task using neural networks involving submodular functions. Our method is compared with two other algorithms focused on optimizing the performance of the worst-case task, and on directly optimizing the performance on the reference distribution itself. We conclude that our novel formulation produces a solution that is locally distributional robust, and computationally inexpensive.
Keywords: distributionally robust optimization (DRO), local distributional robustness, multi-task submodular optimization, relative-entropy (KL) regularization, robust subset selection, greedy submodular maximization, stochastic greedy.
@@ 1 Introduction
Submodular functions have for long been the focus of extensive studies[18, 19, 27, 33], thanks to their propensity to occur naturally in many distinct domains such as economics and algorithmic game theory [10, 17, 50] and machine learning[39, 2, 20, 37, 26, 21, 32, 40, 13, 16], within many distinct problems, ranging from sensor selection [35, 33, 51, 34] to social network modeling [46]. In simple terms, a submodular function is a set function that possesses the diminishing marginal returns property [47]. This property, at first glance, enables an analogy between submodular functions and concave functions in the continuous domain. However, interestingly, a similar analogy may be drawn between submodular functions and convex functions as well, through the use of certain extensions of submodular functions onto the continuous domain [42, 33]. The similarity of submodular functions to both convex and concave functions, the white whales of optimization literature, has also garnered interest in the study of the optimization of submodular functions.
@@
The quintessential submodular optimization problem is that of the maximization of a submodular function under a cardinality constraint. This problem entails selecting a subset out of a ground set $\mathcal{N}$, such that the size of the selection does not exceed a predetermined constraint value, and the value of the selection under a submodular function is as high as possible. The evaluation $f(\mathcal{S})$ of a subset $\mathcal{S}\subseteq\mathcal{N}$ is commonly called the score or utility of the set $\mathcal{S}$. The following formulation, where $f$ is a submodular function, $\mathcal{N}$ is the ground set and $K$ is a positive integer, formalizes this problem[10]:
$\begin{gathered}\max_{\mathcal{S}\subseteq\mathcal{N}}f(\mathcal{S})\\ \text{s.t.}\;\lvert\mathcal{S}\rvert\leq K.\end{gathered}$ (1)
@@
The notion of robustness occurs very frequently in optimization literature and is the object of much consideration. Although there are differing definitions of robustness depending on the context, such as being immune to removals from the solution [8], or to slight changes in parameters or input data[38], and others[7], it always encompasses the general idea of a produced output to a problem staying valid under perturbations to the input. In the context of submodular optimization, one notion of robustness that is commonly encountered is multi-task robustness. As evident by the name, with this notion, the aim is to algorithmically produce a solution set whose utility with respect to a family of $n$ submodular functions $f_{1},\ldots,f_{n}$ is satisfactory.
@@
As it stands, the notion of “satisfactory” is intentionally unspecified, and the problem can be formulated in multiple ways to fit the description. A straightforward interpretation with this goal in mind might lead to what we will call the worst-case formulation [33, 4, 5, 55], which aims to produce a solution that maximizes the worst-performing objective function among all. Using the shorthand notation $n]\vcentcolon=\{1,\ldots,n\}$, this formulation is formalized:
$\begin{gathered}\max_{\mathcal{S}\subseteq\mathcal{N}}\min_{i\in[n]}f_{i}(\mathcal{S})\\ \text{s.t.}\;\lvert\mathcal{S}\rvert\leq K.\end{gathered}$ (2)
@@
A shortcoming of this formulation is that it is too pessimistic[43], dedicating all resources to the maximization of the worst-performing objective function, at the expense of the others. In a scenario where one function is a clear straggler, always scoring significantly lower than all others, this approach may effectively lead to getting low utility on all functions in pursuit of the hopeless aim of maximizing the straggler.
@@
Another naive way to formulate the multi-task robust problem is to adopt what we will call the average-case formulation [43, 52], which aims to directly optimize the mean of all the objective functions. In this case, the formulation would be
@@
$\begin{gathered}\max_{\mathcal{S}\subseteq\mathcal{N}}\frac{1}{n}\sum_{i=1}^{n}f_{i}(\mathcal{S})\\ \text{s.t.}\;\lvert\mathcal{S}\rvert\leq K.\end{gathered}$ (3)
@@
The apparent shortcoming of this formulation is that, contrary to the previous formulation, it provides no guarantees about how bad the utility with respect to any single objective function may be. One or more objective functions could be scoring arbitrarily badly, as long as the other objective functions are scoring well enough to compensate for the underperformers.
@@
The previous two formulations, even with their stated shortcomings, may very well be acceptable approaches, especially in the absence of any additional information. However, we argue that if we possess additional information on the relative importance of the objective functions, we can progress to a more meaningful formulation. Suppose, for instance, that we have access to a reference distribution, a discrete probability distribution $Q\in\Delta_{n}$, where $\Delta_{n}$ indicates the $n$-dimensional simplex. In a practical setting, the reference distribution $Q$ may be obtained by a decision-maker subjectively assigning importance scores to each objective function, or by a frequentist approach, where the value assigned to each function would be inferred from how frequently the task that is modeled by said function is performed by the system. It is reasonable to argue that the reference distribution can facilitate incorporating the intent and priority of the main stakeholders in the problem. The incorporation of this additional information to the previous formulation leads to a generalization of (3), which we can view as the scenario where the reference distribution is assumed to be uniform. This idea is encapsulated in the following formulation:
$\begin{gathered}\max_{\mathcal{S}\subseteq\mathcal{N}}\sum_{i=1}^{n}Q_{i}f_{i}(\mathcal{S})\\ \text{s.t.}\;\lvert\mathcal{S}\rvert\leq K.\end{gathered}$ (4)
@@
We note that this formulation is a also generalization of (2), in that when $Q$ is a one-hot distribution assigning all weight to the worst-performing objective function, this problem becomes equivalent to (2).
@@
Fig. 1: Illustration of the three discrete distributions on the $3$-dimensional simplex[53] that the three discussed approaches optimize. $P_{\text{worst}}$ corresponds to the global worst-case task scenario, assigning a weight of $1$ to the worst-case task and a weight of 0 to all the others, residing on a vertex of the simplex. $P_{\text{avg}}$ assigns uniform weight to all tasks, and lies in the center of the simplex. $Q$ is the reference distribution within a neighborhood of which we want to achieve local robustness. $P^{\ast}$ is the local worst-case distribution within that neighborhood of $Q$.
@@ 1.1 Contribution
Based on this last formulation, we propose our robust formulation, which not only makes use of the reference distribution to simply weigh each objective function but to act as the center of a neighborhood of robustness, within which we aim to be robust. The initial candidate formulation for this is:
$\begin{gathered}\max_{\mathcal{S}\subseteq\mathcal{N}}\min_{P\in\Delta_{n}}\sum^{n}_{i=1}P_{i}f_{i}(\mathcal{S})\\ \text{s.t.}\;\lvert\mathcal{S}\rvert\leq K,\;D(P\;\|\;Q)\leq R,\end{gathered}$ (5)
where the constraint $D(P\;\|\;Q)\leq R$ designates the localization to a specified neighborhood of robustness within which we will maximize our objectives weighed by the worst possible distribution. $D(P\;\|\;Q)$ is intentionally unspecified and could be any statistical distance (although one natural choice quickly arises through analysis). For an investigation of different statistical distances for $D(P\;\|\;Q)$, we refer the reader to Section 3.
Before we proceed with a treatment of (5), we propose one last change to make it more tractable. We work with a Lagrangian relaxation the formulation by introducing the constraint $D(P\;\|\;Q)\leq R$ into the objective via regularization constant $\lambda>0$ rather than strictly enforcing it [9]. With this final addition, our formulation becomes
$\begin{gathered}\max_{\mathcal{S}\subseteq\mathcal{N}}\min_{P\in\Delta_{n}}\sum^{n}_{i=1}P_{i}f_{i}(\mathcal{S})+\lambda D(P\;\|\;Q)\\ \text{s.t.}\;\lvert\mathcal{S}\rvert\leq K.\end{gathered}$ (6)
We propose that this final formulation will be useful in many multi-objective applications where we aim to maintain a performance threshold across all tasks, with respect to a reference distribution indicating the relative importance of each task. In particular, we undertake two such applications in Section 6. The first deals with a low Earth orbit satellite constellation performing multiple tasks involving atmospheric readings and ground coverage, where in critical missions it may be of importance to have satisfactory performance in all objectives. The second involves a task of image summarization, where each objective represents the cosine distance of one image in a given image dataset to the current summary chosen out of the dataset. It is clear that in such a task, one would expect a certain level of performance to be met with respect to all of the images, so that the resulting summary is representative of the whole dataset.
@@ 1.2 Related Works and Significance
Robust submodular optimization. A fundamental work in establishing the notion of robustness in submodular optimization is [35], and the same notion is adopted in the current manuscript. The authors propose an algorithmic solution to the worst-case formulation of (2), namely, the Submodular Saturation Algorithm (SSA). SSA will be discussed considerably in the following sections and will constitute one of the baselines we compare against. In [48, 55], the authors similarly adopt the same formulation of (2) in consideration of robustness. The former generalizes the results of [35], which deals with the cardinality-constrained problem, to cases where knapsack or other matroid constraints are present, including the case where the constraints themselves are submodular. The latter shifts the focus to achieving higher computational efficiency, aiming for a fast and practical algorithm with asymptotic approximation guarantees when the number of objective functions $n$ is increased arbitrarily. In [43], the authors propose a novel notion of robustness, still in the context of multiple objective functions. They first motivate their novel approach by outlining the pessimistic nature of (2). They instead propose an approach that entails the maximization of the $p$th quantile of the objective functions. To simplify, this approach may be viewed as disregarding the worst-performing $p$th of the objectives in favor of maximizing the remaining. Different from all of the previously mentioned works, the present work shifts the focus to the scope of distributional robustness, focusing on localizing our goal of robustness within a subset of $\Delta_{n}$, to a neighborhood of a reference distribution. In [8, 54], the authors adopt one of the other aforementioned notions of robustness. In these works, the notion of robustness is against the post-processing of solutions for the arbitrary or adversarial deletion of a certain number of elements, rather than robustness against multiple submodular objective functions.
@@
Distributionally-robust optimization (DRO). Our work is closely related to and motivated by the DRO paradigm. In a very general sense, DRO is a paradigm where some uncertainty about the nature of the problem is governed by a probability distribution and the goal is to leverage the information we have about this distribution with the goal of optimization or optimal decision-making. In [57], the authors approach a convex optimization problem in the context of DRO. In [49], the scope is broadened to the online optimization of nonconvex functions. In particular, the authors use relative-entropy regularization, which we also employ in this work. The authors of [52], much like our work, deal with distributionally-robust optimization of submodular functions. They work in the presence of stochastic submodular objective functions that are drawn from a given reference distribution and aim to maximize their expected value. They propose an approach that entails a variance-regularized objective, making use of the multilinear extension of submodular functions and the Momentum Frank-Wolfe algorithm. In contrast to this line of work, we leverage DRO for discrete multi-task subset selection and demonstrate the possibility of formulating the distributionally-robust problem as a simple submodular maximization problem. This perspective essentially allows us to produce a distributionally-robust solution at no additional cost to producing a solution to (4).
@@ 2 Preliminaries and Background
We begin by introducing the concept of the marginal return in set functions.
Definition 1 (Marginal return).
Given a set function $f\vcentcolon 2^{\mathcal{N}}\rightarrow\mathbb{R}$ and $\mathcal{S},\mathcal{T}\subseteq\mathcal{N}$, we denote $f(\mathcal{T}\cup\mathcal{S})-f(\mathcal{T})$, the marginal return in $f$ due to adding $\mathcal{S}$ to $\mathcal{T}$, by $\Delta_{f}(\mathcal{S}\mid\mathcal{T})$. When $\mathcal{S}$ is a singleton, i.e., $\mathcal{S}=\{a\}$, we simply write $\Delta_{f}(a\mid\mathcal{T})$.
@@
A set function $f\vcentcolon 2^{\mathcal{N}}\rightarrow\mathbb{R}$ is submodular if for every $\mathcal{S}\subseteq\mathcal{T}\subseteq\mathcal{N}$ and $e\in\mathcal{N}\setminus\mathcal{T},$ it holds that
$\Delta_{f}(e\mid\mathcal{S})\geq\Delta_{f}(e\mid\mathcal{T}).$ (7)
This definition of submodularity highlights the diminishing marginal returns property. The common notion in thinking about a submodular function $f$ is that it scores each set by its utility, as such, we will commonly refer to the value $f(\mathcal{S})$ as the utility, performance, or score of set $\mathcal{S}$.
We further introduce two additional useful properties of set functions, that are sought along submodularity in the derivation of theoretical guarantees for approximation algorithms.
@@
A set function $f\vcentcolon 2^{\mathcal{N}}\rightarrow\mathbb{R}$ is normalized if $f(\emptyset)=0$.
@@
A set function $f\vcentcolon 2^{\mathcal{N}}\rightarrow\mathbb{R}$ is monotone nondecreasing if for every $\mathcal{S}\subseteq\mathcal{T}\subseteq\mathcal{N},$ we have $f(\mathcal{S})\leq f(\mathcal{T}).$
Note that a set function that is both normalized and monotone nondecreasing is necessarily nonnegative.
In practical applications, it is often the case that the objective functions involved are not submodular, but weak submodular. In literature, there are multiple relaxations of submodularity, but for the purposes of this paper, we adopt one that is well-established [58, 11, 23, 25, 30], and reasonable in relaxing the diminishing marginal returns property. To facilitate the definition, we first define the weak-submodularity constant:
@@
$w_{f}\vcentcolon=\max_{(\mathcal{S},\mathcal{T},e)\in\tilde{\mathcal{N}}}\dfrac{\Delta_{f}(e\mid\mathcal{T})}{\Delta_{f}(e\mid\mathcal{S})},$ (8)
where $\tilde{\mathcal{N}}=\{(\mathcal{S},\mathcal{T},e):\mathcal{S}\subseteq\mathcal{T}\subset\mathcal{N},e\in\mathcal{N}\setminus\mathcal{T}\}.$
An intuitive understanding of the WSC is that it is the degree of the maximum violation of the diminishing marginal returns property of a set function. Evidently, a monotone nondecreasing set function $f$ is submodular if and only if its WSC satisfies $w_{f}\leq 1$.
We now define weak submodularity.
@@
A set function $f\vcentcolon 2^{\mathcal{N}}\to\mathbb{R}$ is weak submodular if it is monotone nondecreasing and its WSC satisfies $w_{f}<\infty$.
Possibly the most fundamental result in submodular optimization concerns the maximization of normalized, monotone nondecreasing submodular functions under a cardinality constraint. In this case, the iterative Greedy algorithm, which simply consists of going over the entire set of remaining elements at each iteration and selecting the one with the highest marginal return, enjoys an approximation guarantee of 11/e1-1/e. Although simple in principle, Greedy is demonstrably the optimal approximation algorithm, unless $\text{P}=\text{NP}$.[47]
While Greedy is demonstrably optimal and easy to implement, its naive form is usually avoided due to the computational cost of reevaluating every remaining element at each iteration. Several methods have been proposed to circumvent this shortcoming, two significant ones being Lazy Greedy[44], and Stochastic Greedy[45]. In particular, Stochastic Greedy achieves a reduced computational cost by restricting its evaluations to a subset of the remaining elements sampled uniformly at random at each iteration. Stochastic Greedy also enjoys, in expectation, an approximation ratio similar to that of Greedy, with an additional term that designates the dependence on the cardinality of the randomly sampled subset.
Theorem 1 (Stochastic Greedy approximation ratio[45]).
Let $f\vcentcolon 2^{\mathcal{N}}\rightarrow\mathbb{R}$ be a normalized, monotone nondecreasing submodular function. Let $R=(\lvert\mathcal{N}\rvert/K)\log(1/\epsilon)$ be the size of the sampled set at each iteration of Stochastic Greedy used in the solution of (1), where $\epsilon\in(0,1)$. Then, in expectation, Stochastic Greedy achieves an approximation ratio of $1-1/e-\epsilon$, i.e., it produces a solution $\hat{\mathcal{S}}$ which satisfies
$\mathbb{E}[f(\hat{\mathcal{S}})]\geq(1-1/e-\epsilon)f(\mathcal{S}^{\ast}),$ (9)
where $\mathcal{S}^{*}$ is a solution of (1).
In the formulation of (6), we have intentionally left unspecified the statistical distance $D(P\;\|\;Q)$ in the regularization term. In this section, we will be specifically interested in investigating which choices of $D(\cdot\;\|\;\cdot)$ lead to favorable algorithmic solutions and theoretical guarantees for Problem (6). In particular, statistical distances which preserve the submodular nature of the problem will be of specific interest. For, ensuring that the problem remains submodular allows us to produce a locally robust solution essentially at no additional cost with respect to the non-robust formulation of (1), through familiar tools such as Stochastic Greedy, and enjoying theoretical guarantees such as in Theorem 1.
We first shift our attention to $L^{p}$ metrics. One choice for $D(P\;\|\;Q)$ that leads to a favorable algorithmic solution is the $L^{\infty}$ metric, i.e., $D(P\;\|\;Q)=\lVert P-Q\rVert_{\infty}\vcentcolon=\max_{i\in[n]}\lvert P_{i}-Q_{i}\rvert$. With this choice, (6) becomes
$\begin{gathered}\max_{\mathcal{S}\subseteq\mathcal{N}}\min_{P\in\Delta_{n}}\sum^{n}_{i=1}P_{i}f_{i}(\mathcal{S})+\lambda\lVert P-Q\rVert_{\infty}\\ \text{s.t.}\;\lvert\mathcal{S}\rvert\leq K.\end{gathered}$ (10)
In this case, we can exploit the reduction of the inner maximization of (10) to a linear program to derive a result. Since we know that the optimal point will occur at one of the vertices of the feasible region [6], we have the following equivalence:
$\begin{split}\min_{P\in\Delta_{n}}\sum^{n}_{i=1}P_{i}f_{i}(\mathcal{S})+\lambda\lVert P-Q\rVert_{\infty}=\min_{i\in[n]}f_{i}(\mathcal{S})-\lambda Q_{i}.\end{split}$ (11)
Hence, (10) reduces to
$\begin{gathered}\max_{\mathcal{S}\subseteq\mathcal{N}}\min_{i\in[n]}f_{i}(\mathcal{S})-\lambda Q_{i}\\ \text{s.t.}\;\lvert\mathcal{S}\rvert\leq K.\end{gathered}$ (12)
The main observation about (12) is that it is almost identical the worst-case formulation of (2), however, in addition, the regularization by the reference distribution $Q$ still holds influence. This leads us to propose to use SSA with a slight modification enabling the incorporation of a reference distribution $Q$ in the solution of (12). The full algorithm, which we name Saturate with Preference, is presented in Algorithm 1. It has a function evaluation complexity of $O(\lvert\mathcal{N}\rvert^{2}n\log(n\min_{i\in[n]}f_{i}(\mathcal{N}))$, identical to the standard SSA [35].
It is worth noting that a choice of the $L^{1}$ metric, i.e., $D(P\;\|\;Q)=\lVert P-Q\rVert_{1}$, through the same reduction of the inner problem to a linear program, and the same equivalence of (11), results in the exact same formulation of (12). Since $P$ and $Q$ are discrete probability distributions, one may also view the quantity $\lVert P-Q\rVert_{1}$ as a scaling of the total variation distance, an example of an $f$-divergence, which motivates the discussion of the next section, where we use one specific $f$-divergence with particularly significant theoretical outcomes, namely, the relative entropy or KL-divergence [36].
Algorithm 1 Saturate with Preference
Input: Finite family of monotone nondecreasing submodular functions $f_{1},\ldots f_{n}$, ground set $\mathcal{N}$, integer cardinality bound $0\leq K\leq\lvert\mathcal{N}\rvert$, regularization parameter $\lambda>0$, reference distribution $Q\in\Delta_{n}$
Output: Solution set $\mathcal{S}$